OpenStudy (anonymous):

Help me with this log . log(5-2x) = Log (3x+1)

myininaya (myininaya):

Since we have outsides are the same, just set the insides equal and solve for x.

myininaya (myininaya):

So you did this: 5-2x=3x+1?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Did you have a problem solving that equation?

OpenStudy (anonymous):

do i subtract the 1 from the 5

myininaya (myininaya):

That's a start.

myininaya (myininaya):

So you are saying to do this 4-2x=3x right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Now how do you get your x's together?

OpenStudy (anonymous):

i have no idea

myininaya (myininaya):

Well you want to get your x's together on one side and your non x's together on the side. you have 4-2x on the left and 3x on the right how do we get the 2x on the other side. Well you have minus 2x on the left, to undo subtraction you add.

OpenStudy (anonymous):

so do i add it to 3x ??

myininaya (myininaya):

I was thinking more adding 2x on both sides.

myininaya (myininaya):

adding 3x wouldn't help but it isn't wrong if you do it to both sides but it would just lead to no where

OpenStudy (anonymous):

so how would that look?

myininaya (myininaya):

Take our equation from like a few minutes ago and add 2x on both sides and you show me what you think it looks like.

OpenStudy (anonymous):

2x(4-2x) =3x+2x????

myininaya (myininaya):

put an addition symbol between 2x and (4-2x) since we added 2x on both sides. 2x+(4-2x)+3x+2x

myininaya (myininaya):

Now combine like terms on both sides.

OpenStudy (anonymous):

2x+(4-2x)+5x ??

myininaya (myininaya):

darn it 2x+(4-2x)=3x+2x I made a type-o

myininaya (myininaya):

combine like terms

myininaya (myininaya):

2x+4-2x=?

myininaya (myininaya):

then the other side is 3x+2x=?

OpenStudy (anonymous):

4+1x?? 5x

myininaya (myininaya):

no on the first but yes on second

myininaya (myininaya):

2x+4-2x=4 since 2x-2x=0

myininaya (myininaya):

So we have 2x+(4-2x)=3x+2x can be written as 4 =5x

myininaya (myininaya):

Now to get x by itself we divide both sides by?

OpenStudy (anonymous):

4?